what is the nth term of an arithmetic sequence, what is the 101st term
2,4,6,8,10....

Answer by stanbon(75887)

what is the nth term of an arithmetic sequence, what is the 101st term
2,4,6,8,10....
-----------
a(n)=a(1)=(n-1)d
Your Problem:
a(1)=2
d=4-2=2
a(101)=2+100*2=202
============
Cheers,
Stan H.